Fowler and Woods Enterprises is a publicly traded company that just paid a $2.00 per share dividend. The company is expected to increase its dividend by 25% per year for the next two years. After the second year, the dividend growth rate will be 5% per year for the next two years. After the 4th year, dividends are expected to grow at a constant rate of 3% into the foreseeable future. An analyst estimates that investors in the firm will require a 12% annual return. Based on this information, what is the intrinsic value of the stock today?
Answer #1

D1=(2*1.25)=2.5

D2=(2.5*1.25)=3.125

D3=(3.125*1.05)=3.28125

D4=(3.28125*1.05)=3.4453125

Value after year 4=(D4*Growth rate)/(Required return-Growth rate)

=(3.4453125*1.03)/(0.12-0.03)

=39.4296875

Hence intrinsic value=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.5/1.12+3.125/1.12^2+3.28125/1.12^3+3.4453125/1.12^4+39.4296875/1.12^4

=$34.31(Approx)